Calculate the hardness of a water Sample whose 20 ml required 30 ml EDTA, 10ml of standard


calcium chloride solution, whose strength is equivalent to 300mg of CaCO3 per 200 ml required 20ml


of same EDTA solution.

1
Expert's answer
2021-12-09T14:32:02-0500

The standard water contains 1.5 g of CaCO3/l.

1000 ml of standard hard water (SHW), therefore, contain 15,00 mg of CaCO3.

1 ml of standard water = 15,00/1000 = 1.5 mg of CaCO3 equivalent.

10 ml of EDTA solution = 10 ml of SHW contains

=> 10* 1.5 = 15 mg of CaCO3 equivalent

1ml o EDTA Solution = 15 / 10 = 1.5 mg of CaCo3 equivalent

Calculation of hardness of water:

100 ml of water 30 ml of EDTA solution

30 * 1.5 = 45 mg of CaCO3 equivalent

1000 ml of water = 450 mg of CaCO3 equivalent

Hardness of water = 450 mg / l